Consider an economy described by the following equations.          

Y = C + I + G

 

G = 1, 200              T = 1,000

 

C = 475 + 0.8 (Y − T)  I = 1, 000 − 100 r.

 

MD = 2000 - 8000r

 

MS = 2500

 

i.   Find the equilibrium interest rate and equilibrium output.

 

ii.  Suppose that G decreases to 1, 000, now find the new equilibrium interest rate



and output.



1
Expert's answer
2021-01-21T07:57:45-0500

(a)"AE=Y"

AE=1275-0.8y+2200-100r

"R=Y"

1.8r=3475

34.1

"MD=MS"

=2000-800r=2500

2000-2500=800r

-500=800r

=-0.625%

(B) "MD=MS"

2000-800r=1500

2000-1500=800r

500=800r

=0.625%